OpenStudy (anonymous):

Solve for x. 4(x – b) = x (1 point)

OpenStudy (anonymous):

so which would it be out of these 4: b=4/3x, b=3/4x, x=4/3b, or x=3/4b ?

OpenStudy (anonymous):

Distribute.... 4x - 4b = x 3x = 4b x = 4b/3

OpenStudy (anonymous):

Ohhhh ok thanks!